# Winding quotients for virtual period maps of rank 1

### v1

@@ 1. Introduction
Actually, the pull-back of the function $\mathcal{W}$ to the winding covering space has been studied in quite different contexts in mathematical physics either as the propagator $P_{0}^{\infty}(w_{1},w_{2};\tau,\overline{\tau})$ of BCOV theory by Si Li ([L1] Lemma 3.1, [L2] section 4 and [L-Z] 3.3) and also the propagator $P(z)$ by Dijkgraaf [D] (4.44) on elliptic curves, or as the Bergman kernel function $B(z_{1},z_{2})$ for the case genus =1 in the topological recursion by Eynard [E] p.339 Examples (c.f. also Tyurin [T] p.185 Example. $\omega_{\alpha}$ and Fay [F] (46)). We do not know yet whether such connection of the winding phenomenon with mathmatical physics is common for non-geometric virtual period maps in general or it remains only in these test examples. We need to test further cases [S2, G-S].
@@ 2. The virtual period map $z^{\tau}$ on $\mathbb{C}_{z}^{\times}$
We start to define a virtual period map of rank 1 by the solution of period equation.  [^1] We used in [S2] the term “virtual” for the period integrals whose associated root lattices do not have algebraic geometry origin. We employed in the present note the same term “virtual”. However, we use here the term “virtual” simply in the sense that the exponent $\tau$ of the period equation are non-real complex numbers.
@@ 5. Winding of the domain $\mathbb{C}_{z}^{\times}$ of $z^{\tau}$ to the elliptic curve $E_{\tau}$
where the scaling factor $\frac{\tau}{2\pi\sqrt{-1}}$ comes from the scaling of the group $\Gamma_{g}+\Gamma_{w}=\frac{2\pi\sqrt{-1}}{\tau}(\mathbb{Z}+\mathbb{Z}\tau)$ relative to the lattice $\mathbb{Z}+\mathbb{Z}\tau$ for the $\wp$-function. Thus, the new inversion problem asks to describe the pair coordinates $(\wp(u,\tau),\wp^{\prime}(u,\tau))$ of the point $e\in E_{\tau}$ in terms of $w=\exp(\tau\mathrm{Log}z)$ and $\tau$.  [^2] The realization of $E_{\tau}$ as a plane curve is not unique, but there are severals (e.g. [S2]), and in the present note, we have chosen just one of them, i.e. $\mathrm{A}_{2}$-type realization. It would be also interesting to study the other cases.
@@ 6. $q$-multiplicatively periodic function $\mathcal{W}(w,q)$
In the present section, we study a meromorphic function $\mathcal{W}$ on $\mathbb{C}_{w}^{\times}$ which is invariant by the multiplicative action of $q^{\mathbb{Z}}:=\exp(2\pi\sqrt{-1}\tau\mathbb{Z})$ in order to answer to the new inversion problem.
Following the classical idea of the construction of the $\wp$-function by its partial fractional expantion on the $u$-plane $\mathbb{C}_{u}$ (which is additively doubly periodic, having second order poles at $\mathbb{Z}+\mathbb{Z}\tau$), we construct a meromorphic function $\mathcal{W}$ on $w\in\mathbb{C}_{w}^{\times}$ which is multiplicatively periodic in $w$ by the periods $q^{\mathbb{Z}}$, i.e. $\mathcal{W}(wq,q)=\mathcal{W}(w,q)$, having second order poles on the set $\cup_{m\in\mathbb{Z}}\{1-wq^{m}=0\}$.
@@
whose $m$-th term $\frac{wq^{m}}{(1-wq^{m})^{2}}$ is a rational function having the second order pole at $w=q^{-m}$ which is invariant by the involution: $(w,q)\mapsto(w^{-1},q^{-1})$.
Assertion A. The sum obtained by removing the $m$-th term from $(1)$ converges uniformly on any compact subset of the domain $\mathcal{U}_{m}:=\{(w,q)\in\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}\mid|q|^{-m+1}<|w|<|q|^{-m-1}\}$. So the sum $(1)$ defines a meromorphic function on $\mathcal{U}_{m}$. Since we have the open covering $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}=\cup_{m\in\mathbb{Z}}\ \mathcal{U}_{m}$, the sum $(1)$ converges to a meromorphic function on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$, which we shall denote again by $\mathcal{W}(w,q)$. The function $\mathcal{W}(w,q)$ is $q$-multiplicatively periodic.
$(2)\qquad\qquad\qquad\qquad\qquad\qquad\qquad\mathcal{W}(wq,q)\ \ =\ \ \mathcal{W}(w,q)\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
@@
Actually, the differnce was already calculated by R. Borcherds as follows. In [B] Chapter7, Borcherds give the following Fourier expansion of the $\wp$-function (see also Aoki [A] (3.10)).

### v2

@@ 1. Introduction
Actually, the pull-back of the function $\mathcal{W}$ to the winding covering space has been studied in quite different contexts in mathematical physics either as the propagator $P(z)$ by Dijkgraaf [D] (4.44) on elliptic curves and the propagator $P_{0}^{\infty}(w_{1},w_{2};\tau,\overline{\tau})$ of BCOV theory by Si Li ([L1] Lemma 3.1, [L2] section 4 and [L-Z] 3.3), or as the Bergman kernel function $B(z_{1},z_{2})$ for the case genus =1 in the topological recursion by Eynard [E] p.339 Examples (c.f. also Tyurin [T] p.185 Example. $\omega_{\alpha}$ and Fay [F] (46)). We do not know yet whether such connection of the winding phenomenon with mathmatical physics is common for non-geometric virtual period maps in general or it remains only in these test examples. We need to test further cases [S2, G-S].
@@ 2. The virtual period map $z^{\tau}$ on $\mathbb{C}_{z}^{\times}$
We start to define a virtual period map of rank 1 by the solution of period equation.  [^1] We used in [S2] the term “virtual” for the period integrals whose associated root lattices do not have algebraic geometry origin where the exponents become imaginary numbers. We employ in the present note the term “virtual” only in the sense that the exponent $\tau$ of the period equation is a non-real complex number.
@@ 5. Winding of the domain $\mathbb{C}_{z}^{\times}$ of $z^{\tau}$ to the elliptic curve $E_{\tau}$
where the scaling factor $\frac{\tau}{2\pi\sqrt{-1}}$ comes from the scaling of the group $\Gamma_{g}+\Gamma_{w}=\frac{2\pi\sqrt{-1}}{\tau}(\mathbb{Z}+\mathbb{Z}\tau)$ relative to the lattice $\mathbb{Z}+\mathbb{Z}\tau$ for the $\wp$-function. Thus, the new inversion problem asks to describe the pair coordinates $(\wp(u,\tau),\wp^{\prime}(u,\tau))$ of the point $e\in E_{\tau}$ in terms of $w=\exp(\tau\mathrm{Log}z)$ and $\tau$.  [^2] The realization of $E_{\tau}$ as a plane curve is not unique, but there are severals (e.g. [S2]). In the present note, we have chosen $\mathrm{A}_{2}$-type realization. It would be also interesting to study the other cases. $\mathrm{B}_{2}$-type and $\mathrm{G}_{2}$-type
@@ 6. $q$-multiplicatively periodic function $\mathcal{W}(w,q)$
In the present section, we study a meromorphic function $\mathcal{W}$ on $\mathbb{C}_{w}^{\times}$ which is invariant by the multiplicative action of $q^{\mathbb{Z}}\!:=\!\exp(2\pi\sqrt{-1}\tau\mathbb{Z})$ on $\mathbb{C}_{w}^{\times}$ in order to answer to the new inversion problem.
Following the classical idea of the construction of the $\wp$-function by its partial fractional expansion on the $u$-plane $\mathbb{C}_{u}$ (which is additively doubly periodic, having second order poles at $\mathbb{Z}+\mathbb{Z}\tau$), we construct a meromorphic function $\mathcal{W}$ by its partial fractional expansion on $w\in\mathbb{C}_{w}^{\times}$ which is multiplicatively periodic by the periods $q^{\mathbb{Z}}$, $\mathcal{W}(wq,q)=\mathcal{W}(w,q)$ and is involutive $\mathcal{W}(w^{-1},q^{-1})=\mathcal{W}(w,q)$, having second order poles on the set $\cup_{m\in\mathbb{Z}}\{1-wq^{m}=0\}$.
@@
where the $m$-th term $\frac{wq^{m}}{(1-wq^{m})^{2}}$ is a rational function having the second order pole at $w=q^{-m}$ which is invariant by the involution: $(w,q)\mapsto(w^{-1},q^{-1})$.
Assertion A. For any $m\in\mathbb{Z}$, the sum obtained by removing the $m$-th term from $(1)$ converges uniformly on any compact subset of the domain $\mathcal{U}_{m}:=\{(w,q)\in\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}\mid|q|^{-m+1}<|w|<|q|^{-m-1}\}$. So the sum $(1)$ defines a meromorphic function on $\mathcal{U}_{m}$.
Since we have the open covering $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}=\cup_{m\in\mathbb{Z}}\ \mathcal{U}_{m}$, the sum $(1)$ converges to a meromorphic function on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$, which we shall denote again by $\mathcal{W}(w,q)$ which is $q$-multiplicatively periodic and involution invariant. [^3] The two properties together with the second order poles characterize $\mathcal{W}(w,q)$ up to an involutive function in $q$.
$(2)\qquad\qquad\qquad\qquad\mathcal{W}(wq,q)\ \ =\ \ \mathcal{W}(w,q)\qquad\text{and}\qquad\mathcal{W}(w^{-1},q^{-1})\ \ =\ \ \mathcal{W}(w,q).\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
@@
Actually, the differnce was already known by R. Borcherds as follows. In [B] Chapter7, Borcherds give the following Fourier expansion of the $\wp$-function (see also Aoki [A] (3.10), Lang [L] Chap.4 Sec.2 Prop. 2-3.).